[figure: 9.7 cm]

FERROMAGNETIC ORDERING IN UFe_{0.5}Ni_{0.5}Al

V.H. Tran^{a}, R. Troć^{a}, G. André^{b} and F. Bourée^{b}

aInstitute of Low Temperature and Structure Research, Polish Academy of Sciences,

50-950 Wroław, Poland

bLaboratoire Léon Brillouin, (CEA-CNRS), CEA/Saclay, 91191 Gif-Sur-Yvette Cedex, France

The solid solutions between non-magnetic UFeAl and antiferromagnetic UNiAl quite unexpectedly yield ferromagnetic properties of the UFe_{1−x}Ni_{x}Al alloys around x = 0.5 [Troć et al. PRB 51 (1995) 3003]. In this work we investigate the nature of magnetic ordering of UFe_{0.5}Ni_{0.5}Al by means of neutron powder diffraction and specific heat measurements. The investigated alloy crystallizes in the hexagonal Fe_{2}P-type structure, and the transition a ferromagnetic state occurs at T_{C} = 55 K, as previously reported. The magnetic structure consists of the ferromagnetic (001)-planes with the magnetic moments parallel to the c-axis. At 1.5 K, the ordered uranium moment (μ_{ord}) is about 0.6 μ_{B}. Surprisingly, no clear λ-type anomaly in the specific heat (C_{p}) is observed at T_{C}. The temperature dependence of C_{p} may be described by an expression involving an energy gap in the magnon spectrum. The electronic coefficient of the specific heat is modestly enhanced and amounts to about 47 mJ/mole-U K^{2}. This feature together with a small value of μ_{ord} point to the strongly hybridized character of the 5f-d magnetism in UFe_{0.5}Ni_{0.5}Al.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : oral

Corresponding author : V.H. Tran

Address for correspondence : Inst. Niskich Temp. i Badan Strukt. PAN 50-950 Wroclaw

Email address : vhtran@int.pan.wroc.pl